# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v1

@@ VI-B ERADS Quantum Flux Ultra Refined
The main differences with RADEA are the use of a memory factor to guide the mutation in certain directions and an adaptive $F$ and $CR$ strategy. It is unclear, why the LLM generated the ”Quantum Flux” component in the algorithm’s name, but it can be assumed that LLM hallucinations [44] will also occur when applied to code generation tasks.
It is interesting to observe that both RADEA and ERADS are in principle very similar, with the key difference that ERADS uses adaptive parameter control and a memory factor for mutation, while RADEA uses an archive of best solutions.
VI-C Performance Analysis in Higher Dimensions
The two best-found algorithms using the LLaMEA framework are evaluated against the most relevant state-of-the-art optimizers, namely the previously used baselines BIPOP-CMA-ES, CMA-ES and Differential Evolution (DE), and the JADE algorithm [43], all of which are using recommended hyper-parameter settings. The baselines all originate from the IOH analyzer benchmark data set [45]. In Figure 9, the empirical attainment function of the proposed algorithms and baselines is shown for dimension $d\in\{10,20\}$ using results from all BBOB functions, $5$ instances per function and $5$ random seeds ($25$ runs per BBOB function in total). It is interesting to observe that while ERADS performs well in $5$ dimensional problems, it fails to generalize well to higher dimensions, while RADEA has very good performance in higher dimensions as well (in terms of anytime performance, which is the area below these curves). ERADS performs overall slightly better than JADE and standard DE and as good as the CMA-ES and BIPOP-CMA-ES.
See Appendix II for the detailed convergence curves per BBOB function in $5d$, $10d$ and $20d$.
@@ VI-C Performance Analysis in Higher Dimensions
Figure 9: The empirical attainment function (EAF) estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. EAF for the best two algorithm (RADEA and ERADS) and the baselines; CMA-ES, DE, BIPOP-CMA-ES and JADE, averaged over all 24 BBOB functions in (from top to bottom ) $10d$ and $20d$ respectively.

### v2

@@ VI Analysis of Best Algorithms
The main differences with DE are the use of a memory factor to guide the mutation in certain directions and an adaptive $F$ and $CR$ strategy. It is unclear, why the LLM generated the ”Quantum Flux” component in the algorithm’s name, but it can be assumed that LLM hallucinations [43] will also occur when applied to code generation tasks.
The proposed ERADS algorithm seems most similar to the existing JADE [44] algorithm from literature. There are, however, some key differences which are, to the best of our knowledge, novel and have not been published before.
The key differences primarily revolve around the use of a memory vector and a memory factor in combination with parameter adaptation. JADE uses an optional archive with the best solutions which primarily serves to maintain diversity. Furthermore, JADE also features adaptive control parameters (F and CR) that adjust dynamically based on the success rates of previous generations, aiming to optimise these parameters in real time to improve performance. It is interesting to observe that most of the best algorithms proposed by the LLM are similar to Differential Evolution (or Hybrids of CMA and DE), this could indicate a bias towards this kind of algorithms in the LLM, or it could indicate that overall these kind of algorithms work well for the specific benchmark suite.
VI-A Performance Analysis in Higher Dimensions
The best-found algorithms using the LLaMEA framework are evaluated against the most relevant state-of-the-art optimizers, namely the previously used baselines CMA-best, CMA-ES and Differential Evolution (DE), CMA-ES and DE are using recommended hyper-parameter settings. The baselines all originate from the IOH analyzer benchmark data set [45]. In Figure 9, the empirical attainment function of the proposed algorithms and baselines is shown for dimension $d\in\{10,20\}$ using results from all BBOB functions, $5$ instances per function and $5$ random seeds ($25$ runs per BBOB function in total). It is interesting to observe that while ERADS performs well in $5$ dimensional problems, it fails to generalize well to higher dimensions, while EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism both have very good performance in higher dimensions for the first $2000$ function evaluations (even better than CMA-best), but around $10000$ function evaluations CMA-best and CMA-ES outperform the other algorithms clearly. When looking at convergence curves per function, we note that this is mainly due to the very hard functions in the Benchmark ($23$ and $24$), See Appendix II for the detailed convergence curves per BBOB function in $5d$, $10d$ and $20d$.
@@ VI-A Performance Analysis in Higher Dimensions
Figure 9: The empirical attainment function (EAF) estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. EAF for the best algorithm per configuration and the baselines; CMA-ES, CMA-best and DE, averaged over all 24 BBOB functions in (from top to bottom) $10d$ and $20d$ respectively.